# A Unique Bosonic Symmetry in a 4D Field-Theoretic System

### v1

@@ 3 Algebraic Structures: Discrete Duality and Continuous Nilpotent Symmetry Operators
$\bar{B}_{\mu\nu}\to\pm\,\bar{\cal B}_{\mu\nu},\quad\bar{\cal B}_{\mu\nu}\to\mp\,\bar{B}_{\mu\nu},\quad F_{\mu}\to\pm\,\bar{f}_{\mu},\quad\bar{f}_{\mu}\to\mp\,F_{\mu}.$ (16)
Within the framework of BRST formalism, there is an alternative physical realization of the well-known mathematical relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives $(\delta)d$ of differential geometry in terms of (i) the off-shell nilpotent (i.e. $s_{ab}^{2}=0,\;s_{ad}^{2}=0$), infinitesimal and continuous anti-BRST symmetry transformations $s_{ab}$ [cf. Eq. (10)] and the anti-co-BRST symmetry transformations $s_{ad}$ [cf. Eq. (9)], and (ii) the discrete dualityWe christen the discrete symmetry transformations in (13) and (16) as the duality symmetry transformations because, under these discrete symmetry transformations, we note that (i) the gauge-fixing term of the Abelian 1-form basic gauge field transforms to the kinetic term for the Abelian 3-form basic gauge field and vice-versa, and (ii) the gauge-fixing term for the basic Abelian 3-form gauge field exchanges with the kinetic term for the Abelian 1-form basic gauge field. As stated earlier, the kinetic terms and gauge-fixing terms for the basic Abelian 3-form and 1-form gauge fields owe their mathematical origins to the exterior and dual-exterior derivatives of differential geometry, respectively. symmetry transformations (13) and (16). This statement can be mathematically expressed in the following explicit form, namely;
$s_{ab}\,\Sigma$ $=$ $-\,*\,s_{ad}\,*\,\Sigma,\qquad\Sigma=\beta_{\mu},\;\bar{\beta}_{\mu},\,B_{4},\,B_{5},\qquad s_{ab}\,\widetilde{\Phi}=+\,*\,s_{ad}\,*\,\widetilde{\Phi}$,
$\widetilde{\Phi}$ $=$ $A_{\mu\nu\sigma},\,B_{\mu\nu},\,{\cal B}_{\mu\nu},\,\bar{B}_{\mu\nu},\,\bar{\cal B}_{\mu\nu},\,\bar{C}_{\mu\nu},\,C_{\mu\nu},\,A_{\mu},\,\phi_{\mu},\widetilde{\phi}_{\mu},\bar{f}_{\mu},\,F_{\mu},\,f_{\mu},\,\bar{F}_{\mu},$ (17)
@@
where the $\mp$ signs in the relationships $s_{ab}\,\Sigma=-\,*\,s_{ad}\,*\,\Sigma$ and $s_{ab}\,\widetilde{\Phi}=+\,*\,s_{ad}\,*\,\widetilde{\Phi}$ are just the opposite of our observations in (15) with the replacement: $\Phi\to\widetilde{\Phi}$. .
@@
turn out to be zero provided we use the three specific CF-type restrictions (i.e. ${\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) which define the subspace of qauntum fields where the co-BRST and anti-co-BRST symmetry transformation operators absolutely anticommute (i.e. $\{s_{d},\;s_{ad}\}=0$). The absolute anticommutativity (i.e. $\{s_{d},\;s_{ad}\}=0$) is satisfied trivially for the rest the quantum fields of our 4D BRST-quantized theory. Hence, we finally conclude that the absolute anticommutativity (i.e. $\{s_{b},\;s_{ab}\}=0$ and $\{s_{d},\;s_{ad}\}=0$) between the off-shell nilpotent (i.e. $s_{(a)b}^{2}=0,\,s_{(a)d}^{2}=0$) (anti-)BRST ($s_{(a)b}$) and (anti-)dual-BRST ($s_{(a)d}$) symmetry transformation operators are satisfied provided we utilize the specific set of only three CF-type restrictions (out of the total four).
@@
Furthermore, we would like to add that the anticommutators in (20) do not define a new set of bosonic symmetry transformations because the Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ do not (i) remain invariant under (20), and/or (ii) transform to the total spacetime derivatives under (20). It is worthwhile to mention that the anticommutators (20) are true for the rest of the fields of our theory [except the ones that are present in (21)]. It is very interesting to point out, in our present context, that if we make the choice: $B_{2}=B_{3}=B_{4}=B_{5}=0$, we find that the anticommutators in (20) exactly become zero. However, even under this limit, the off-shell nilpotent symmetry transformations $s_{(a)b}$ and $s_{(a)d}$ as well as the bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ exist. Second, the following non-trivial anticmmmutators between the off-shell nilpotent symmetry operators, namely;
@@
$s_{d}\bar{f}_{\mu}=-\,*\,s_{b}\,*\bar{f}_{\mu}=\partial_{\mu}B_{5},\qquad s_{d}F_{\mu}=-\,*\,s_{b}\,*\,F_{\mu}=-\,\partial_{\mu}B_{3},$
@@
Finally, it is the peculiarity of our 4D BRST-quantized theory that the reciprocal relationships of (14) and (17) also exist and they are found to be true. In other words, it can be verified clearly that (i) the relationships $s_{b}\,\Sigma=+\,*\,s_{d}\,*\,\Sigma$ and $s_{b}\,\Phi=-\,*\,s_{d}\,*\,\Phi$, and (ii) the relationships $s_{ad}\,\Sigma=-\,*\,s_{ab}\,*\,\Sigma$ and $s_{ad}\,\Phi=+\,*\,s_{ab}\,*\,\Phi,$ are also true. However, the actual mapping between the nilpotent symmetry and cohomological operators $(\delta)d$ would be dictated by the ghost number considerations (cf. our section five for more discussions).
4 Bosonic Symmetry Transformations: Uniqueness Property and the Full Set of CF-Type Restrictions
The non-nilpotent (i.e. $s_{\omega}^{2}\neq 0$) bosonic symmetry operator $s_{\omega}$ [cf. Eq. (22)] can operate on the individual fields of this Lagrangian density ${\cal L}_{(B)}$. The following infinitesimal, continuous and non-nilpotent symmetry transformations under $s_{\omega}=\{s_{b},\;s_{d}\}$, namely;
@@ 4 Bosonic Symmetry Transformations: Uniqueness Property and the Full Set of CF-Type Restrictions
$s_{\omega}A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}B^{\sigma\rho}+\partial_{\mu}B_{1},\qquad s_{\omega}\bar{\beta}_{\mu}=+\,\partial_{\mu}B_{5},\qquad s_{\omega}\beta_{\mu}=+\,\partial_{\mu}B_{4},$
@@
lead to the transformation of the Lagrangian density ${\cal L}_{(B)}$ as follows
@@
thereby rendering the action integral $S=\int d^{4}x\,{\cal L}_{(B)}$ invariant (i.e. $s_{\omega}S=0$). The transformation in (25) can be easily derived by using our observations in equations (5) and (6). In other words, we have $s_{\omega}{\cal L}_{(B)}=\big(s_{b}\,s_{d}+s_{d}\,s_{b}\big)\,{\cal L}_{(B)}$ which implies the transformation (25)
@@
$s_{\bar{\omega}}A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}\bar{B}^{\sigma\rho}-\partial_{\mu}B_{1},\qquad s_{\bar{\omega}}\bar{\beta}_{\mu}=-\,\partial_{\mu}B_{5},\qquad s_{\bar{\omega}}\beta_{\mu}=-\,\partial_{\mu}B_{4},$
@@
Taking into account the sum of transformations (24) and (26), we obtain the following:
@@
It is crystal clear that the relationship $s_{\omega}\,+\,s_{\bar{\omega}}=0$ is true provided we consider our whole BRST-quantized theory on the space of quantum fields where all the four CF-type restrictions (i.e. $B_{\mu\nu}+\bar{B}_{\mu\nu}=(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}),\;{\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) are satisfied together. This observation should be contrasted against the backdrop of our observations in the context of the proof of the absolute anticommutativity (i) between the off-shell nilpotent BRST and anti-BRST symmetry transformation operators [cf. Eq. (18)], and (ii) between the off-shell nilpotent co-BRST and anti-co-BRST symmetry transformation operators [cf. Eq. (19)], where only three (out of the total four) specific types of CF-type restrictions have been invoked. We conclude, ultimately, that there is a unique bosonic symmetry transformation operator (e.g. $s_{\omega}$) in our theory.
@@
$-\,\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}-B_{4}\,\partial^{\mu}B_{5}+B_{5}\,\partial^{\mu}B_{4}$

### v2

@@ 3 Algebraic Structures: Discrete Duality and Continuous Nilpotent Symmetry Operators
$\bar{B}_{\mu\nu}\to\pm\,i\,\bar{\cal B}_{\mu\nu},\quad\bar{\cal B}_{\mu\nu}\to\pm\,i\,\bar{B}_{\mu\nu},\quad F_{\mu}\to\pm\,i\,\bar{f}_{\mu},\quad\bar{f}_{\mu}\to\pm\,i\,F_{\mu}.$ (16)
Within the framework of BRST formalism, there is an alternative physical realization of the well-known mathematical relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives $(\delta)d$ of differential geometry in terms of (i) the off-shell nilpotent (i.e. $s_{ab}^{2}=0,\;s_{ad}^{2}=0$), infinitesimal and continuous anti-BRST symmetry transformations $s_{ab}$ [cf. Eq. (10)] and the anti-co-BRST symmetry transformations $s_{ad}$ [cf. Eq. (9)], and (ii) the discrete duality******We christen the discrete symmetry transformations in (13) and (16) as the duality symmetry transformations because, under these discrete symmetry transformations, we note that (i) the gauge-fixing term of the Abelian 1-form basic gauge field transforms to the kinetic term for the Abelian 3-form basic gauge field and vice-versa, and (ii) the gauge-fixing term for the basic Abelian 3-form gauge field exchanges with the kinetic term for the Abelian 1-form basic gauge field. As stated earlier, the kinetic terms and gauge-fixing terms for the basic Abelian 3-form and 1-form gauge fields owe their mathematical origins to the exterior and dual-exterior derivatives of differential geometry, respectively. symmetry transformations (13) and (16). This statement can be mathematically expressed in the following explicit form, namely;
$s_{ad}\,\Sigma$ $=$ $+\,*\,s_{ab}\,*\,\Sigma,\qquad\Sigma=\beta_{\mu},\;\bar{\beta}_{\mu},\,B_{4},\,B_{5},\qquad s_{ad}\,\widetilde{\Phi}=-\,*\,s_{ab}\,*\,\widetilde{\Phi}$,
$\widetilde{\Phi}$ $=$ $A_{\mu\nu\sigma},\,\bar{B}_{\mu\nu},\,\bar{\cal B}_{\mu\nu},\,\bar{C}_{\mu\nu},\,C_{\mu\nu},\,A_{\mu},\,\phi_{\mu},\widetilde{\phi}_{\mu},\,f_{\mu},\,\bar{F}_{\mu},$ (17)
@@
where the $\pm$ signs in the relationships $s_{ad}\,\Sigma=+\,*\,s_{ab}\,*\,\Sigma$ and $s_{ad}\,\widetilde{\Phi}=+\,*\,s_{ab}\,*\,\widetilde{\Phi}$ are exactly the same as our observations in (15) with the replacement: $\Phi\to\widetilde{\Phi}$. The mathematical relationships in equations (14) and (17) merely establish that the off-shell nilpotent (i.e. $s_{(a)b}^{2}=0,\;s_{(a)d}^{2}=0$) (anti-)BRST $s_{(a)b}$ and (anti-)co-BRST $s_{(a)d}$ symmetry transformation operators are dual to one-another (cf. Sec. 5 for more comments).
@@
turn out to be zero provided we use the three specific CF-type restrictions (i.e. ${\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) which define the subspace of qauntum fields where the co-BRST and anti-co-BRST symmetry transformation operators absolutely anticommute (i.e. $\{s_{d},\;s_{ad}\}=0$). The absolute anticommutativity (i.e. $\{s_{d},\;s_{ad}\}=0$) is satisfied trivially for the rest the quantum fields of our 4D BRST-quantized theory. Hence, we finally conclude that the absolute anticommutativity (i.e. $\{s_{b},\;s_{ab}\}=0$ and $\{s_{d},\;s_{ad}\}=0$) between the off-shell nilpotent (i.e. $s_{(a)b}^{2}=0,\,s_{(a)d}^{2}=0$) (anti-)BRST ($s_{(a)b}$) and (anti-)dual-BRST ($s_{(a)d}$) symmetry transformation operators are satisfied provided we utilize the specific set of only††††††It has been shown in our earlier works [16,17] that the requirements of the simultaneous (anti-)BRST and (anti-)co-BRST invariance(s) of the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ also lead to the derivations of only three specific sets of the CF-type restrictions (out of the total four). three CF-type restrictions (out of the total four).
@@
Furthermore, we would like to add that the anticommutators in (20) do not define a new set of bosonic symmetry transformations because, even though the Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ transform to the total spacetime derivatives under (20), they change the ghost number of the specific fields (i.e. $\phi_{\mu},\,\widetilde{\phi}_{\mu},\,\beta_{\mu},\,\bar{\beta}_{\mu}$) by ± $\pm\,2$ (cf. Appendix A for more details). It is worthwhile to mention that the anticommutators (20) are true for the rest of the fields of our theory [except the ones that are present in (21)]. It is very interesting to point out, in our present context, that if we make the choice: $B_{2}=B_{3}=B_{4}=B_{5}=0$, we find that the anticommutators in (20) exactly become zero. However, even under this limit, the off-shell nilpotent symmetry transformations $s_{(a)b}$ and $s_{(a)d}$ as well as the bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ exist [cf. Eqs. (24),(26) below]. Second, the following non-trivial anticmmmutators between the nilpotent symmetry operators, namely;
@@
$s_{d}\bar{f}_{\mu}=-\,*\,s_{b}\,*\bar{f}_{\mu}=\partial_{\mu}B_{5},\qquad s_{d}F_{\mu}=-\,*\,s_{b}\,*\,F_{\mu}=+\,\partial_{\mu}B_{3},$
@@
Finally, it is worthwhile to point out that, even though the mathematical relationships in equations (14) and (17) are similar, it is the ghost number considerations‡‡‡‡‡‡It is straightforward to note that the pair of nilpotent operators ($s_{d},\;s_{ab}$) lower the ghost number of a field on which they operate. On the contrary, the the pair of nilpotent operators ($s_{b},\;s_{ad}$) raise ghost number of a field by one (on which theory operate). These observations are similar to the consequences that emerge out when the nilpotent (co-)exterior derivatives operate on a given differential form. that decide, out of the pairs ($s_{d},\;s_{ad}$) and ($s_{b},\;s_{ab}$), which operator corresponds to the nilpotent (i.e. $\delta^{2}=0$) co-exterior derivative $\delta$ and which operator represents the nilpotent (i.e. $d^{2}=0$) co-exterior derivative $d$ of differential geometry (cf. Sec. 5 for more discussions).
4 Bosonic Symmetry Transformations: Uniqueness Property and Full Set of CF-Type Restrictions
The non-nilpotent (i.e. $s_{\omega}^{2}\neq 0$) bosonic symmetry operator $s_{\omega}$ [cf. Eq. (22)] can operate on the individual fields of this Lagrangian density ${\cal L}_{(B)}$. The following infinitesimal, continuous and non-nilpotent symmetry transformations (under $s_{\omega}=\{s_{b},\;s_{d}\}$), namely;
@@ 4 Bosonic Symmetry Transformations: Uniqueness Property and Full Set of CF-Type Restrictions
$s_{\omega}A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}B^{\sigma\rho}-\partial_{\mu}B_{1},\qquad s_{\omega}\beta_{\mu}=+\,\partial_{\mu}B_{4},\qquad s_{\omega}\bar{\beta}_{\mu}=+\,\partial_{\mu}B_{5},$
@@
lead to the transformation of the Lagrangian density ${\cal L}_{(B)}$ as
@@
thereby rendering the action integral $S=\int d^{4}x\,{\cal L}_{(B)}$ invariant (i.e. $s_{\omega}S=0$). The transformation in (25) can be easily derived by using our observations in equations (5) and (6). In other words, we have $s_{\omega}{\cal L}_{(B)}=\big(s_{b}\,s_{d}+s_{d}\,s_{b}\big)\,{\cal L}_{(B)}$ which implies the transformation (25).
@@
$s_{\bar{\omega}}A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}\bar{B}^{\sigma\rho}+\partial_{\mu}B_{1},\qquad s_{\bar{\omega}}\beta_{\mu}=-\,\partial_{\mu}B_{4},\qquad s_{\bar{\omega}}\bar{\beta}_{\mu}=-\,\partial_{\mu}B_{5},$
@@
Taking into account the sum of transformations (24) and (26), we obtain:
@@
It is crystal clear that the operator relationship $s_{\omega}\,+\,s_{\bar{\omega}}=0$ is true provided we consider our whole BRST-quantized theory on the subspace of quantum fields where all the four CF-type restrictions (i.e. $B_{\mu\nu}+\bar{B}_{\mu\nu}=(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}),\;{\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) are satisfied together. This observation should be contrasted against the backdrop of our observations in the context of the proof of the absolute anticommutativity (i) between the nilpotent BRST and anti-BRST symmetry transformation operators [cf. Eq. (18)], and (ii) between the nilpotent co-BRST and anti-co-BRST symmetry transformation operators [cf. Eq. (19)], where only three (out of the total four) specific types of CF-type restrictions have been invoked. We conclude, ultimately, that there is a unique bosonic symmetry transformation operator (e.g. $s_{\omega}$) in our theory.
@@
$-\,\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}+B_{5}\,\partial^{\mu}B_{4}-B_{4}\,\partial^{\mu}B_{5}$